The midpoint between Nancy and Moe's houses is (-1,3). Nancy's home is at (x,-1) and Moe's house is located at (2,y).
1. Find the x-y coordinate where Nancy's home is located.
2. Find the x-y coordinate where Moe's home is located.
3. Find the distance between Nancy and Moe's homes.
Help is appreciated! :)

Answer by lwsshak3(11628)

The midpoint between Nancy and Moe's houses is (-1,3). Nancy's home is at (x,-1) and Moe's house is located at (2,y).
1. Find the x-y coordinate where Nancy's home is located.
2. Find the x-y coordinate where Moe's home is located.
3. Find the distance between Nancy and Moe's homes.
***
midpoint formula: (x1+x2)/2, (y1+y2)/2
(x+2)/2, (-1+y)/2=(-1,3)
..
(x+2)/2=-1
x+2=-2
x=-4
..
(-1+y)/2=3
-1+y=6
y=7
1. (x-y) coordinate where Nancy's home is located: (-4,-1)
2. (x-y) coordinate where Moe's home is located: (2,7)
..
distance formula:√[(x1-x2)^2+(y1-y2)^2]
√[(-4-2)^2+(-1-7)^2)]=√(36+64)=√100=10
3. distance between Nancy and Moe's homes=10